OpenStudy (anonymous):

The perimeter of a kite is 45 meters. If one side measures 10 meters find the measure of the opposite side.

OpenStudy (anonymous):

PerimeterKite=2x+2y We know P=45 and one side let's say x=10 45=2*10+2y 45=20+2y 25=2y y=12.5m